OpenStudy (anonymous):

The slope is -2 and the line passes through (4,-5). what is the equation?

OpenStudy (anonymous):

y=-2x+3

OpenStudy (anonymous):

Not really. These things are the hardest for me.

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

first you have to identify this as a direct line, in general we write it as y=mx+c where m is the slope and c is the intercept. m=-2 is given.apply the point (4,-5) which will give you the c.

OpenStudy (anonymous):

y=-2x+3 where do the 4 and -5 go?

OpenStudy (anonymous):

since you know the slope (m=-2) apply it, y=-2x+c to get the c apply (4,-5),that is x=4 and y=-5

OpenStudy (anonymous):

Oh. Ok.

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Wait, im confused.

OpenStudy (anonymous):

so that would be -5=-2(4)+c ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do you get that to y=2x+3 ?

OpenStudy (anonymous):

-5=-2(4)+c solve this you yeild c=3 so now you have everything for your direct line, y=-2x+3

OpenStudy (anonymous):

so my next problem is Slope= 2/5 and line passes through (2,1/2)

OpenStudy (anonymous):

so -1/2=2/5(2)+c ??

OpenStudy (anonymous):

same as above,yes you are correct but y should be 1/2.

OpenStudy (anonymous):

sorry i wrote it wrong. (2,-1/2)

OpenStudy (anonymous):

then you applied it correctly :-)

OpenStudy (anonymous):

now im just confused with solving. especially with fractions

OpenStudy (anonymous):

why don't you use a calculator?

OpenStudy (anonymous):

-13/10 ?